The csmemo Class for L^{A}TEX 2ε

Martin Reddy

April 15, 1997

1 Introduction

This document describes the csmemo class (v1.1) which has been written for L^{A}TEX 2ε. This class provides a generic internal memorandum sheet for use in the department of Computer Science. It modifies the macro \maketitle to produce a memo page, and provides a number of other macros to customise the content of the memo. You can use this class by simply starting your L^{A}TEX file with:

\documentclass{csmemo}

2 Specifying the Memo Headers

The \maketitle command produces a memo page. The following commands can be used before \maketitle in order to customise the headers of the memorandum.

• \MemoTo{<name(s)>} : the name(s) of the intended recipient(s). If you have multiple names, then you can separate each with a \newline command in order to place each on a new line. (You should always specify this field.)

• \MemoFrom{<name>} : the sender of the memo. (You should always specify this field.)

• \MemoCC{<name(s)>} : the name(s) of any people who are to receive a carbon copy of the memo. That is, the memo is not directly addressed to these recipients, but you want them to see it. (This field is optional.)

• \MemoDate{<date>} : the date that you wish to appear on the memo. If you do not specify this, then the current date will be used. (This field is optional.)

• \MemoSubject{<subject>} : the subject matter of the memo. (You should always specify this field.)

• \MemoSpecial{<descr>}{<text>} : if you want to add a line to the memo headers that is not one of To, From, CC, Date, or Subject, then you can use this command to specify any descriptor and text. This will appear after the From line, or the CC line if present. (This field is optional.)

3 Customising the Address

In addition to the above commands, the csmemo class allows you to perform some customisation of the address information which appears to the top right of the memo. Accordingly, the following (optional) commands are available:

• \AddressName : the name of the sender of the memo. This will appear at the top of the address (directly below the Department of Computer Science logo).

• \AddressTitle : the title of the sender of the memo, e.g. Senior Lecturer, or Head of Department.

• \AddressAltTitle : an alternative title that the sender may hold, e.g. Head of Planning Unit.

• \AddressExt : the 4-digit telephone extension for the sender of the memo. By default, the ‘Tel:’ line will contain the general enquiries contact number. However, if you specify this field, then you can include your personal office number.

• \AddressEmail : this will allow you to display the sender’s e-mail address. This is presented after the address and telephone details.

4 Example L^{A}TEX 2ε Code

\documentclass{csmemo}
\begin{document}
\AddressName{Mr Frodo Baggins}
\AddressTitle{Head Hobbit of Hobbiton}
\AddressEmail{F.Baggins@bagend.shire.middle-earth}
\AddressExt{0000}
\MemoTo{Samwise Gamgee\newline Meriadoc Brandybuck \newline Peregrim Took}
\MemoFrom{Frodo Baggins}
\MemoCC{Gandalf the Grey}
\MemoDate{September 22, 3018 (Third Age)}
\MemoSubject{The Journey to Rivendell}
\maketitle
Just to remind you all that we leave tomorrow. Remember to bring lots
of warm clothing, and don't forget the hogsweed!
\end{document}

 Above L^{A}TEX source available on-line in /usr/local/share/TeX/tex/latex/examples